A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC.
1
Expert's answer
2020-05-05T18:43:24-0400

Assume that the carbon resistor has a resistance of 500Ω at "T_0 = 0\degree C" and the temperature coefficient of resistance for carbon is "\alpha = -0.0048 \Omega/\degree C" . Then, according to the following formula:

"R(T) = R(T_0)(1-\alpha (T-T_0))"

the resistance at "T = 50\degree C" will be:

"R(50\degree C) = 500\Omega(1--0.0048 \Omega/\degree C(50\degree C-0\degree C)) = 380 \Omega" .


Answer. 380 Ohms.